OpenStudy (anonymous):

Solve the system of equations using substitution. y=2x-10 y=4x-8

OpenStudy (anonymous):

First solve for x(or y) in one of the equations Since we already know y=4x-8, lets put that into the first equation y=2x-10 (4x-8)=2x-10 Solve for x 2x=-2 x=-1 Now plug -1 in for x on one of the other equations y=2(-1)-10 y=-2-10=-12 So we know that x=-1 and y=-12